# Heterogeneity-Aware Cooperative Federated Edge Learning with Adaptive Computation and Communication Compression

### v1

@@ 1 Introduction
As a distributed machine learning (ML) paradigm that has gained significant attention recently, federated learning (FL) allows edge devices to collaboratively learn a shared model while keeping their training data locally, which offers several advantages such as privacy protection and better efficiency compared to centralized ML paradigm [1]. When deployed over mobile edge networks, the standard FL framework is cloud-based and relies on a central cloud server to coordinate the learning process and aggregate model updates from all edge devices. However, due to the long-distance and limited-bandwidth transmissions between an edge device and the remote cloud, the model training in cloud-based FL is inevitably slow and fails to meet the latency requirements of delay-sensitive intelligent applications [2]. Moreover, the reliance on a central server presents the risk of a single point of failure in cloud-based FL.
With the increasing deployment of computation and storage resources at the edge in 5G-and-beyond networks, an edge-based FL framework called cooperative federated edge learning (CFEL) is emerging as a promising alternative to cloud-based FL [3, 4, 5, 6, 7]. In this framework, edge servers co-located with the mobile base stations are responsible for the coordination among their proximate edge devices. Edge servers only communicate with their neighboring servers to avoid the communication bottleneck at the central cloud. Moreover, by eliminating the need for a single server to collect local model updates from all devices, CFEL can effectively mitigate the risk of a single point of failure and provide a scalable FL framework for vast numbers of edge devices distributed across a wide area.
Despite its immense potential, CFEL faces several challenges that hinder training efficiency: 1) System Heterogeneity. In the CFEL system, the edge devices are notably heterogeneous, displaying a wide range of capabilities [8, 9]. For example, there can be substantial differences in computing power, such as CPU frequency and battery life, with some edge devices being up to ten times more powerful than others. Similarly, the communication capabilities, including bandwidth and throughput, also vary significantly among these edge devices. 2) Dynamic State. Since the environment of participating devices may fluctuate over time, the available resources (e.g., CPU states, channel states, and battery life) for training on edge devices could be dynamic [10, 11]. Due to system heterogeneity and dynamic state, the computation and communication resources may vary significantly across all devices, leading to the straggler problem during the training with large training delay and high energy consumption. How to achieve efficient model training under system heterogeneity and dynamic state in CFEL remains largely unknown.
@@ 2 Related works
FL over mobile edge networks encounters significant challenges such as high training latency and energy consumption. To improve communication efficiency, research efforts have been made to reduce the size of communication data via model compression [13, 14, 15, 16, 17, 18, 19]. An orthogonal research direction is to optimize the local update frequency [20, 21] based on the observation that when the local update frequency increase, the number of communication rounds between devices and the central server may be reduced. Moreover, some recent studies [12, 11] jointly optimize local updating frequency and model compression ratio to speed up training and reduce resource consumption. However, all aforementioned works [20, 21, 13, 14, 15, 16, 17, 18, 19, 12, 11] assume a central server that aggregates model updates from all edge devices and hence are not directly applicable to the CFEL system considered in this paper.
@@ 3.1 CFEL: Cooperative Federated Edge Learning
As shown in Fig. 1, we consider a CFEL system consisting of $N$ devices that are distributed among $m$ clusters. Each cluster $i\in[m]$ contains one edge server and a set of devices $\mathcal{S}_{i}$ with $N_{i}=|\mathcal{S}_{i}|$. A device $n\in\mathcal{S}_{i}$ is associated with edge server $i$ by some predefined criteria such as physical distance and wireless network protocols. The set of all edge devices in the system is denoted by $\mathcal{S}=\cup_{i=1}^{m}\mathcal{S}_{i}$ with $N=|\mathcal{S}|$. The edge backhaul network facilitates communication between edge servers. The topology of the edge backhaul network is represented as a connected and undirected graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of edge servers and $\mathcal{E}$ is the set of communication links between edge servers. Some main notations used in the paper is summarized in Table I.
@@
The classic FedAvg algorithm cannot be directly applied to the CFEL system due to the presence of multiple clusters and lack of a central server for model aggregation. To solve (1) under the CFEL network architecture, new learning algorithms (e.g., MLL-SGD [3], SD-FEEL [4], P-FedAvg [5] and CE-FedAvg [6]) have been proposed. These algorithms mainly comprise three key stages as follows. (i) Local model update: each device performs multiple local iterations to update its local model via the mini-batch SGD; (ii) Intra-cluster aggregation: after updating the local model, each device uploads its updated local model to the corresponding edge server, and the edge server computes the averaged edger server model and sends it back to the associated devices; and (iii) Inter-cluster aggregation: after multiple rounds of intra-cluster aggregations, each edge server communicates with its neighboring edge servers to do the inter-cluster aggregation, and sends the intra-cluster aggregated model to its associated devices. Although prior work has demonstrated the potential of such a new learning algorithm, how to address challenges such as system heterogeneity and dynamic state in CFEL remains unknown.
@@ 3.2 Heterogeneity-Aware CFEL
5:           Each edge device $n$ uploads its local parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, and $\nu_{n}^{l,r}$ to the coordinator by Algorithm 2.
6:           Coordinator finds the optimal compression ratios $\theta_{n}^{l,r}$ and probabilities $\rho_{n}^{l,r}$ for $n\in\mathcal{S}_{i}$ according to Algorithm 3 and sends them to $\mathcal{S}_{i}$.
7:           for each device $n\in\mathcal{S}_{i}$ in parallel do
8:              $\mathbf{x}_{n}^{l,r,0}\leftarrow\mathbf{y}_{i}^{l,r}$
9:              for each local iteration $s=0,\ldots,\tau-1$ do
10:                 Compute a mini-batch gradient $\mathbf{g}_{n}^{l,r,s}$ with probability $\rho_{n}^{l,r}$.
11:                 $\mathbf{x}_{n}^{l,r,s+1}\leftarrow\mathbf{x}_{n}^{l,r,s}-\eta\mathbf{g}_{n}^{l,r,s}$
12:              end for
13:              $\Delta_{n}^{l,r}\leftarrow Q(\mathbf{x}_{n}^{l,r,\tau}-\mathbf{x}_{n}^{l,r,0})$
14:           end for
15:           $\mathbf{y}_{i}^{l,r+1}\leftarrow\mathbf{y}_{i}^{l,r}+\frac{1}{N_{i}}\sum_{n\in\mathcal{S}_{i}}\Delta_{n}^{l,r}$
16:        end for
17:        $\mathbf{y}_{i}^{l+1,0}\leftarrow\sum_{j\in\{i\}\cup\mathcal{N}_{i}}\mathbf{H}_{j,i}\mathbf{y}_{j}^{l,q}$
18:     end for
@@
To address the aforementioned challenges in CFEL, we present a new scheme called HCEF, which allows adaptive control of local update frequency and gradient compression. To monitor the dynamic state and heterogenous devices’ computing and communication capabilities during the training, similar to prior decentralized FL works [22, 23, 9, 24], we assume that there exists a coordinator and ignore the cost (e.g., bandwidth consumption and time cost) for information collection since the size of this information (e.g., 100-300KB [25]) is much smaller than that of model parameters. Furthermore, the role of the coordinator can be dynamically assigned to different edge servers. This means that if one coordinator fails, other edge servers can quickly take over the coordination tasks, ensuring that the training process continues without interruption. We summarize the HCEF scheme for CFEL in Algorithm 1. The overall training process of HCEF is divided into $\phi$ global rounds wherein each cluster first performs $q$ edge rounds of intra-cluster collaboration independently and then communicates with other clusters for inter-cluster collaboration.
Specifically, at the beginning of the $r$-th edge round in the $l$-th global round, denoted as the $(r,l)$-th round, each device needs to upload a range of parameters to the coordinator, including the training parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, and communication parameter $\nu_{n}^{l,r}$ (line 5). This process will be elaborated in Section 5. Then, the coordinator uses Algorithm 3 to determine the compression ratio $\theta_{n}^{l,r}$ and local update probability $\rho_{n}^{l,r}$ for each device by the uploaded training parameters. This procedure will be elaborated in Section 5.
Next, the coordinator sends $\theta_{n}^{l,r}$ and $\rho_{n}^{l,r}$ to each device $n$ (line 6). Meanwhile, each edge server $i$ broadcasts its current model $\mathbf{y}_{i}^{l,r}$ to the associated devices $n\in\mathcal{S}_{i}$. Let $\mathbf{x}_{n}^{l,r,s}$ denote the local model of device $n$ at the $s$-th local iteration of $(r,l)$-th round. Each device $n$ updates its local model $\mathbf{x}_{n}^{l,r,0}$ to the corresponding current edge model (line 8). It is important to note that, considering the heterogeneous computation capabilities of the devices, each device $n$ performs its updates independently and potentially asynchronously. The updating rule for each device can be described by the following:
@@
where $\eta$ is the local learning rate, and $\mathbf{g}_{n}^{l,r,s}$ is a random stochastic gradient determined by the adaptive local update probability (line 11).
After $\tau$ local iterations, each device compresses the model updates using the compression operator $Q(\cdot)$ according to the compression ratios $\theta_{n}^{l,r}=k_{n}^{l,r}/d$ where $k_{n}^{l,r}$ is the number of uploaded parameters for device $n$ at the $(r,l)-$th round (line 13) and sends the compressed updates $\Delta_{n}^{l,r}$ to the associated edge server $i$ for intra-cluster aggregation (line 15). Specifically, once receiving all compressed model updates from the associated devices, each edge server $i$ updates its edge model $\mathbf{y}_{i}^{l,r}$ by aggregating the compressed model updates. It’s crucial to recognize that the compression ratio $\theta_{n}^{l,r}=k_{n}^{l,r}/d$ varies across devices, influenced by factors such as each device’s bandwidth resources, communication capabilities, and the quality of the model being updated. Higher compression ratios, while requiring more bandwidth resources and extending communication time, can enhance model quality and increase training accuracy. Conversely, lower ratios deteriorate model quality but conserve bandwidth resources and improve communication efficiency. Therefore, selecting the optimal compression ratio requires a careful balance, considering the specific bandwidth limitations, communication needs, and model quality requirements of each device.
After $q$ edge rounds, the edge servers engage in inter-cluster aggregation by communicating with their neighboring servers via the edge backhaul (line 17). Specifically, each edge server updates its model by taking the average with neighboring servers using gossip protocol as follows:
@@ 3.3 Problem Formulation
Time Consumption Model. We consider both computing and communication time. Note that the download time from the edge server to each device is ignored because the download bandwidth is typically much larger than the upload bandwidth in device-to-edge communications [2]. We define $\mu_{n}^{l,r}$ and $\nu_{n}^{l,r}$ as the computing time for one local iteration and upload time of one full model for device $n$ at $(r,l)$-th round, respectively, and $\mathcal{T}_{i,i^{\prime}}^{l}$ as the transmission time from cluster $i$ to its neighbor $i^{\prime}$. Taking into account the heterogeneous local update probability $\rho_{n}^{l,r}$ and compression ratio $\theta_{n}^{l,r}$, the total expected training time for $\phi$ global rounds can be determined as:
@@
Here, the averaged model is widely used in the convergence analysis of decentralized algorithms in the literature [30, 4, 6]. In practice, the edge models will reach consensus after convergence, and the averaged model will become the single global model.
With the above time and energy models, our problem can be formulated as the following constrained optimization problem:
@@ 4 Convergence Analysis
    If $(i,j)\in E$, then $\mathbf{H}_{i,j}>0$; otherwise, $\mathbf{H}_{i,j}=0$.
@@
    The magnitudes of all eigenvalues except the largest one are strictly less than 1, i.e., $\zeta=\max\{{|\lambda_{2}(\mathbf{H})|,|\lambda_{m}(\mathbf{H})|}\}<\lambda_{1}(\mathbf{H})=1$
@@
$+\frac{4\eta L}{NT}\sum_{t=0}^{T-1}\sum_{n=1}^{N}(2-\theta_{n}^{t})\rho_{n}^{t}\sigma^{2}+\frac{4(F(\mathbf{x}^{1})-F_{\inf})}{\eta T}$
@@
$\frac{1}{\phi q}$ $\sum_{l=0}^{\phi-1}\sum_{r=0}^{q-1}\mathbb{E}\|\nabla F(\mathbf{u}^{l,r})\|^{2}\leq\frac{4(F(\mathbf{x}^{0,0,1})-F_{\inf})}{\eta\phi q\tau}$
@@
The first term is identical to centralized SGD [31]. As $pq\tau\rightarrow\infty$, this term goes to zero. The second term is the stochastic error related to the compression ratios and local update probabilities. If the stochastic gradients are highly diverged, the local update frequency is high, or the compression ratio is low, then the stochastic error will be more significant. In line with [3], we observe that when the updating probability is low, the stochastic error becomes smaller. The third term reveals the additive error depends on the local update probability. If there is no adaptive local update frequency, i.e., $\rho_{k}^{l,r}=1$, the error is equal to zero. Furthermore, Theorem 1 shows that the iteration complexity of HCEF is $\mathcal{O}(1/(\eta\phi q\tau)+\eta\sigma^{2}+(\eta+1)G^{2})$.
@@
When there is neither compression nor adaptive local updates, meaning that $\theta_{n}^{l,r}=\rho_{n}^{l,r}=1$ for all $n,l,r$, and all devices communicate with a single edge server with $q=1$, our HCEF algorithm reduces to the FedAvg algorithm. In this case, if the learning rate satisfies $\eta={1}/{(L\sqrt{T})}$ when $T>\tau^{4}$, the iteration complexity of HCEF satisfies $\mathcal{O}({1}/{\sqrt{T}})+\mathcal{O}({(\sigma^{2}+G^{2})}/{\sqrt{T}}).$ This coincides with the complexity of FedAvg given in [35, 36].

### v3

@@ 1 Introduction
As a distributed machine learning (ML) paradigm that has gained significant attention recently, federated learning (FL) allows edge devices to collaboratively learn a shared model while keeping their training data locally, which offers several advantages such as privacy protection and better efficiency compared to centralized ML paradigm [1]. When deployed over mobile edge networks [2, 3, 4], the standard FL framework is cloud-based and relies on a central cloud server to coordinate the learning process and aggregate model updates from all edge devices. However, due to the long-distance and limited-bandwidth transmissions between an edge device and the remote cloud, the model training in cloud-based FL is inevitably slow and fails to meet the latency requirements of delay-sensitive intelligent applications [5]. Moreover, the reliance on a central server presents the risk of a single point of failure in cloud-based FL.
With the increasing deployment of computation and storage resources at the edge in 5G-and-beyond networks, an edge-based FL framework called cooperative federated edge learning (CFEL) is emerging as a promising alternative to cloud-based FL [6, 7, 8, 9]. In this framework, edge servers co-located with the mobile base stations are responsible for the coordination among their proximate edge devices. Edge servers only communicate with their neighboring servers to avoid the communication bottleneck at the central cloud. Moreover, by eliminating the need for a single server to collect local model updates from all devices, CFEL can effectively mitigate the risk of a single point of failure and provide a scalable FL framework for vast numbers of edge devices distributed across a wide area.
Despite its immense potential, CFEL faces several challenges that hinder training efficiency: 1) System Heterogeneity. In the CFEL system, the edge devices are notably heterogeneous, displaying a wide range of capabilities. For example, there can be substantial differences in computing power, such as CPU frequency and battery life, with some edge devices being up to ten times more powerful than others. Similarly, the communication capabilities, including bandwidth and throughput, also vary significantly among these edge devices. 2) Dynamic State. Since the environment of participating devices may fluctuate over time, the available resources (e.g., CPU states, channel states, and battery life) for training on edge devices could be dynamic. Due to system heterogeneity and dynamic state, the computation and communication resources may vary significantly across all devices, leading to the straggler problem during the training with large training delay and high energy consumption. How to achieve efficient model training under system heterogeneity and dynamic state in CFEL remains largely unknown.
@@ 2 Related works
FL over mobile edge networks encounters significant challenges such as high training latency and energy consumption. To improve communication efficiency, research efforts have been made to reduce the size of communication data via model compression [12, 13, 14, 15, 16, 17, 18, 19]. An orthogonal research direction is to optimize the local update frequency [20, 21] based on the observation that when the local update frequency increase, the number of communication rounds between devices and the central server may be reduced. Moreover, some recent studies [10, 11] jointly optimize local updating frequency and model compression ratio to speed up training and reduce resource consumption. However, Nori et al. [10] applied the identical local update frequency and model compression ratio to all heterogeneous devices, which limits the full utilization of each device’s capacity. Xu et al. [11] used an adaptive local update frequency and model compression ratio for heterogeneous devices, but their approach, like the aforementioned works [20, 21, 12, 13, 14, 15, 16, 19, 17, 18, 10, 11], assume a central server that aggregates model updates from all edge devices and hence are not directly applicable to the CFEL system considered in this paper.
@@ 3.1 CFEL: Cooperative Federated Edge Learning
As shown in Fig. 1, we consider a CFEL system consisting of $N$ devices that are distributed among $m$ clusters. Each cluster $i\in[m]$ contains one edge server and a set of devices $\mathcal{S}_{i}$ with $N_{i}=|\mathcal{S}_{i}|$. A device $n\in\mathcal{S}_{i}$ is associated with edge server $i$ by some predefined criteria such as physical distance and wireless network protocols. The set of all edge devices in the system is denoted by $\mathcal{S}=\cup_{i=1}^{m}\mathcal{S}_{i}$ with $N=|\mathcal{S}|$. The edge backhaul network facilitates communication between edge servers. The topology of the edge backhaul network is represented as a connected and undirected graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of edge servers and $\mathcal{E}$ is the set of communication links between edge servers. Some main notations used in the paper are summarized in Table I.
@@
The classic FedAvg algorithm cannot be directly applied to the CFEL system due to the presence of multiple clusters and lack of a central server for model aggregation. To solve (1) under the CFEL network architecture, new learning algorithms (e.g., MLL-SGD [6], SD-FEEL [7], P-FedAvg [8] and CE-FedAvg [9]) have been proposed. These algorithms mainly comprise three key stages as follows. (i) Local model update: each device performs multiple local iterations to update its local model via the mini-batch SGD; (ii) Intra-cluster aggregation: after updating the local model, each device uploads its updated local model to the corresponding edge server, and the edge server computes the averaged edger server model and send it back to the associated devices; and (iii) Inter-cluster aggregation: after multiple rounds of intra-cluster aggregations, each edge server communicates with its neighboring edge servers to do the inter-cluster aggregation, and sends the intra-cluster aggregated model to its associated devices. Although prior work has demonstrated the potential of such a new learning algorithm, how to address challenges such as system heterogeneity and dynamic state in CFEL remains unknown.
@@ 3.2 Heterogeneity-Aware CFEL
5:           Each edge server $i$ sends $\mathbf{y}_{i}^{l,r}$ to its associated devices $n\in\mathcal{S}_{i}$.
6:           Each edge device $n$ uploads its local parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, and $\nu_{n}^{l,r}$ to the coordinator by Algorithm 2.
7:           Coordinator finds the optimal compression ratios $\theta_{n}^{l,r}$ and probabilities $\rho_{n}^{l,r}$ for $n\in\mathcal{S}_{i}$ according to Algorithm 3 and sends them to $\mathcal{S}_{i}$.
8:           for each device $n\in\mathcal{S}_{i}$ in parallel do
9:              $\mathbf{x}_{n}^{l,r,0}\leftarrow\mathbf{y}_{i}^{l,r}$
10:              for each local iteration $s=0,\ldots,\tau-1$ do
11:                 Compute a mini-batch gradient $\mathbf{g}_{n}^{l,r,s}$ with probability $\rho_{n}^{l,r}$.
12:                 $\mathbf{x}_{n}^{l,r,s+1}\leftarrow\mathbf{x}_{n}^{l,r,s}-\eta\mathbf{g}_{n}^{l,r,s}$
13:              end for
14:              $\Delta_{n}^{l,r}\leftarrow Q(\mathbf{x}_{n}^{l,r,\tau}-\mathbf{x}_{n}^{l,r,0})$
15:           end for
16:           $\mathbf{y}_{i}^{l,r+1}\leftarrow\mathbf{y}_{i}^{l,r}+\frac{1}{N_{i}}\sum_{n\in\mathcal{S}_{i}}\Delta_{n}^{l,r}$
17:        end for
18:        $\mathbf{y}_{i}^{l+1,0}\leftarrow\sum_{j\in\{i\}\cup\mathcal{N}_{i}}\mathbf{H}_{j,i}\mathbf{y}_{j}^{l,q}$
@@
20:  end for
To address the aforementioned challenges in CFEL, we present a new scheme called HCEF, which allows adaptive control of local update frequency and gradient compression. To monitor the dynamic state and heterogenous devices’ computing and communication capabilities during the training, similar to prior decentralized FL works [22, 23, 24], we assume that there exists a coordinator and ignore the cost (e.g., bandwidth consumption and time cost) for information collection since the size of this information (e.g., 100-300KB [25]) is much smaller than that of model parameters. Furthermore, the role of the coordinator can be dynamically assigned to different edge servers. This means that if one coordinator fails, other edge servers can quickly take over the coordination tasks, ensuring that the training process continues without interruption. We summarize the HCEF scheme for CFEL in Algorithm 1. The overall training process of HCEF is divided into $\phi$ global rounds wherein each cluster first performs $q$ edge rounds of intra-cluster collaboration independently and then communicates with other clusters for inter-cluster collaboration.
Specifically, at the beginning of the $r$-th edge round in the $l$-th global round, denoted as the $(r,l)$-th round, each edge server $i$ broadcasts its edge model $\mathbf{y}_{i}^{l,r}$ to the associated devices $n\in\mathcal{S}_{i}$ (line 5). After that, each device $n$ needs to estimate and upload several parameters to the coordinator, including the training parameters $(\sigma_{n}^{l,r})^{2}$, $(G_{n}^{l,r})^{2}$, computing parameters $\mu_{n}^{l,r}$, $\alpha_{n}^{l,r}$, and communication parameter $\nu_{n}^{l,r}$ (line 6). This process will be elaborated in Section 5.2. Then, the coordinator uses Algorithm 3 to determine the compression ratio $\theta_{n}^{l,r}$ and local update probability $\rho_{n}^{l,r}$ for each device by the uploaded training parameters. This procedure will be elaborated in Section 5.
Next, considering the heterogeneous computation capabilities of the devices, each device $n$ performs its updates independently and potentially asynchronously. The updating rule for each device can be described by the following:
@@
where $\eta$ is the local learning rate, and $\mathbf{g}_{n}^{l,r,s}$ is a random stochastic gradient determined by the adaptive local update probability (line 11-12).
After $\tau$ local iterations, each device compresses the model updates using the compression operator $Q(\cdot)$ according to the compression ratios $\theta_{n}^{l,r}=k_{n}^{l,r}/d$ where $k_{n}^{l,r}$ is the number of uploaded parameters for device $n$ at the $(r,l)-$th round (line 14) and sends the compressed updates $\Delta_{n}^{l,r}$ to the associated edge server $i$ for intra-cluster aggregation (line 16). Specifically, once receiving all compressed model updates from the associated devices, each edge server $i$ updates its edge model $\mathbf{y}_{i}^{l,r}$ by aggregating the compressed model updates. It’s crucial to recognize that the compression ratio $\theta_{n}^{l,r}=k_{n}^{l,r}/d$ varies across devices, influenced by factors such as each device’s bandwidth resources, communication capabilities, and the quality of the model being updated. Higher compression ratios, while requiring more bandwidth resources and extending communication time, can enhance model quality and increase training accuracy. Conversely, lower ratios deteriorate model quality but conserve bandwidth resources and improve communication efficiency. Therefore, selecting the optimal compression ratio requires a careful balance, considering the specific bandwidth limitations, communication needs, and model quality requirements of each device.
After $q$ edge rounds, the edge servers engage in inter-cluster aggregation by communicating with their neighboring servers via the edge backhaul (line 18). Specifically, each edge server updates its model by taking the average with neighboring servers using gossip protocol as follows:
@@ 3.3 Problem Formulation
Time Consumption Model. We consider both computation and communication time. Since the download bandwidth is typically much higher than the upload bandwidth [5], we only focus on the uploading time in the system modeling without loss of generality as in [11, 29, 23, 9]. We define $\mu_{n}^{l,r}$ and $\nu_{n}^{l,r}$ as the computing time for one local iteration and upload time of one full model for device $n$ at $(r,l)$-th round, respectively, and $\mathcal{T}_{i,i^{\prime}}^{l}$ as the transmission time from cluster $i$ to its neighbor $i^{\prime}$. Taking into account the heterogeneous local update probability $\rho_{n}^{l,r}$ and compression ratio $\theta_{n}^{l,r}$, the total expected training time for $\phi$ global rounds can be determined as:
@@
Here, the averaged model is widely used in the convergence analysis of decentralized algorithms in the literature [31, 7, 9]. In practice, the edge models will reach consensus after convergence, and the averaged model will become the single global model. With the above time and energy models, our problem can be formulated as the following constrained optimization problem:
@@ 4 Convergence Analysis
    $(i,j)\in\mathcal{E}$, then $\mathbf{H}_{i,j}>0$; otherwise, $\mathbf{H}_{i,j}=0$.
@@
    The magnitudes of all eigenvalues except the largest one are strictly less than 1, i.e., $\zeta=\max\{{|\lambda_{2}(\mathbf{H})|,|\lambda_{m}(\mathbf{H})|}\}<\lambda_{1}(\mathbf{H})=1$.
@@
$+\frac{4\eta L}{NT}\sum_{t=0}^{T-1}\sum_{n=1}^{N}(2-\theta_{n}^{t})\rho_{n}^{t}\sigma^{2}+\frac{4(F(\mathbf{u}^{0})-F_{\inf})}{\eta T}$
@@
$\frac{1}{\phi q}$ $\sum_{l=0}^{\phi-1}\sum_{r=0}^{q-1}\mathbb{E}\|\nabla F(\mathbf{u}^{l,r})\|^{2}\leq\frac{4(F(\mathbf{u}^{0,0})-F_{\inf})}{\eta\phi q\tau}$
@@
The first term is identical to centralized SGD [32]. As $\phi q\tau\rightarrow\infty$, this term goes to zero. The second term is the stochastic error related to the compression ratios and local update probabilities. If the stochastic gradients are highly diverged, the local update frequency is high, or the compression ratio is low, then the stochastic error will be more significant. In line with [6], we observe that when the updating probability is low, the stochastic error becomes smaller. The third term reveals the additive error depends on the local update probability. If there is no adaptive local update frequency, i.e., $\rho_{k}^{l,r}=1$, the error is equal to zero. Furthermore, Theorem 1 shows that the iteration complexity of HCEF is $\mathcal{O}(1/(\eta\phi q\tau)+\eta\sigma^{2}+(\eta+1)G^{2})$.
@@
When there is neither compression nor adaptive local updates, meaning that $\theta_{n}^{l,r}=\rho_{n}^{l,r}=1$ for all $n,l,r$, and all devices communicate with a single edge server with $q=1$, our HCEF algorithm reduces to the FedAvg algorithm. In this case, if the learning rate satisfies $\eta={1}/{(L\sqrt{T})}$ when $T>\tau^{4}$, the iteration complexity of HCEF satisfies $\mathcal{O}({1}/{\sqrt{T}})+\mathcal{O}({(\sigma^{2}+G^{2})}/{\sqrt{T}}).$ This coincides with the complexity of FedAvg given in [36].